OpenStudy (anonymous):

Evaluate the expression when b= -4 b2+5b+7

OpenStudy (anonymous):

all you have to do is substitute in -4 for b -4(2)+5(-4)+7 -8-20+7=-21

OpenStudy (anonymous):

so what u say is the answer is -8-20+7=21 alright

OpenStudy (anonymous):

-4(2)+5(-4)+7= -8+-21+7= -29+7+=-22

OpenStudy (anonymous):

foxy 5(-4)=-20 not -21

OpenStudy (anonymous):

oops yea its -20

OpenStudy (anonymous):

so the answer its -21

OpenStudy (anonymous):

yes, i double checked my math after you answered just to make sure @niman make sure you have the - in front of 21 for your answer